OpenStudy (anonymous):

Each side of the square measures 8sqrt(5) centimeters, and the area of the square is 32x square centimeters. What is the value of x? A. 5 B. 10 C. 20 D. 40 i think its b????

OpenStudy (anonymous):

@StudyGurl14 could you help me again sorry i feel like a pain:D

OpenStudy (anonymous):

@Zarkon could you help me??

OpenStudy (anonymous):

@zepdrix could you help me?

zepdrix (zepdrix):

Hey there c:

OpenStudy (anonymous):

hi :D

zepdrix (zepdrix):

\[\Large\rm A=\ell^2\]Since we're dealing with a square, the length is the same as the width, so our area is given by length times length, or length squared.

zepdrix (zepdrix):

Each side length is given by \(\Large\rm \ell =8\sqrt5\) and the area is given to be \(\Large\rm 32x\).

zepdrix (zepdrix):

So we have:\[\Large\rm A=(\ell)^2\]\[\Large\rm 32x=\left(8\sqrt5\right)^2\]

zepdrix (zepdrix):

\[\Large\rm 32x=64\cdot 5\]\[\Large\rm 32\cdot x=32\cdot10\]You came up with B?

zepdrix (zepdrix):

Yayyyy good job \c:/

OpenStudy (anonymous):

thankyou sooooooo much!!! :D :D